• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 08:02:14 on June17,2010

    Entry number 327043

    keyword=owl shift summary

    First 3 hours we had no beam.

    Note, R-HRS dipole is not working.

    03:00 ... Do the beam recovery.

    03:40 ... Ramp up to 50 uA on Pb/D tgt 3

    04:00 ... At 50 uA. Doing production, L-HRS only.

    Production Runs 4620 - 4626

    (run 4621 is junk, 4620 is a ramp up)

    Beam tuning at end of shift.